OpenStudy (anonymous):

HELP ASAP! MEDAL+FAN Simplify into one fraction -2 / x + 3 - 4/x-5

jagr2713 (jagr2713):

\[\frac{ -2 }{ x+3 }\frac{ -4 }{ x-5 }\] Like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@jager2713 that is correct, could you help me?

jagr2713 (jagr2713):

what do u think we do first? an is the an answer choice

OpenStudy (anonymous):

I think we add the top numbers first

OpenStudy (anonymous):

the answers are -6x - 2 / x+3)(x-5 6(x-1)/ (x+3)(x-5) -6/(x+3)(x-5) 2/(x+3)(x-5) I think it's the last one

OpenStudy (anonymous):

@jager2713 are you still there?

jagr2713 (jagr2713):

the bottom is correct but you messed up at the top Ok we simplify 2/x+3 X -4/x-5 to get \[\frac{ 2 }{ x+3 }\frac{ 4 }{ x-5 }\]

jagr2713 (jagr2713):

Now what do we do?

OpenStudy (anonymous):

oh wait, the problem is really -2/x+3 minus 4/x-5

jagr2713 (jagr2713):

No we multiply 2*4=8 so we have \[\frac{ 8 }{ (x+3)(x-5) }\]

jagr2713 (jagr2713):

So whats the answer :D

OpenStudy (anonymous):

but i thought we add -2 and 4

OpenStudy (anonymous):

because we are suppose to subtract the fractions

jagr2713 (jagr2713):

It was until we simplify and got positive 2 and 4

OpenStudy (anonymous):

oh okay so that answer we got was 8/x+3, x-5, what else is there to do?

jagr2713 (jagr2713):

Correct

jagr2713 (jagr2713):

Thats it and dont forget the parentheses \[\frac{ 8 }{ (x+3)(x-5) }\]

OpenStudy (anonymous):

but that is not one of the answer choices that i gave you

jagr2713 (jagr2713):

Hmm i see

OpenStudy (anonymous):

i'll do what i think is right, but could you see if you could help me figure this one out? (-5w^10 + 10w^8+5w^6) / (5w^5)

OpenStudy (anonymous):

the answers are -w^5 + 5w^3 = w -5w^5 + 10w^8 + w -w^5 +10^8+ 5w^6 and -w^5 + 2w^3 + w